OpenStudy (karringtonlewis):

A day care program has an average daily expense of $75.00. The standard deviation is $15.00. The owner takes a sample of 64 bills. What is the probability the mean of his sample will be between $70.00 and $80.00?

OpenStudy (loser66):

Central theorem.

OpenStudy (loser66):

mean \(\mu =75\), standard deviation = 15, n =64 \(\bar X\) is sample mean \(P(70< \bar X <80)\) = ? Apply Central theorem \(P(\dfrac{70-\mu}{\sigma/\sqrt n}<\dfrac{\bar X-\mu }{\sigma/\sqrt n}<\dfrac{80-\mu}{\sigma/\sqrt n})=?\)

OpenStudy (loser66):

the middle term is normal distribution \(N\sim (75,15)\) you just put all in calculator.